OpenStudy (anonymous):

(5 * 3-1) * (50/5^2) @JoStudy

OpenStudy (anonymous):

OK \[(5*3-1)*(\frac{ 50 }{ 5^{2} })\]

OpenStudy (anonymous):

\(\huge((5*3-1)*(\frac{ 50 }{ 5^{2} })\)

OpenStudy (anonymous):

Remember exponents so parenthesis first. exponents would be the answer here

OpenStudy (anonymous):

Actully if there are no parethesis around 3-1 then she's wrong

OpenStudy (anonymous):

i was thinking the same

OpenStudy (anonymous):

ok so let's go ahead and start solving :) solve the exponent first :)

OpenStudy (anonymous):

it would be 50/5^2 ??

OpenStudy (anonymous):

50/25

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

\(\huge((5*3-1)*(\frac{ 50 }{ 25})\)

OpenStudy (anonymous):

28 ??

OpenStudy (anonymous):

YEP!

OpenStudy (anonymous):

\(\huge\color{royalblue} {AWESOME ~NICOLE!}\)